In memory address, Hexadecimal what is 0x1Fe0 and 0x2Fe1

Also how to read it? what's its binary equivalent?

How can an assembly program take value from it ? with an example.
Answer #1

Short for hexadecimal. A number system based on 16 values (called base 16), which is explained in this appendix. Uses the 16 numerals 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, A, B, C, D, E, and F. Hex numbers are often followed by a lowercase h to indicate they are in hex.

Memory Addressing :-

Computers often display memory addresses in the hex number system. You must either “think in hex” or convert to decimal. It’s really easier, with a little practice, to think in hex. Here’s the way it works: Memory addresses are displayed as two hex numbers. An example is C800:5. The part to the left of the colon (C800) is called the segment address, and the part to the right of the colon (5) is called the offset. The offset value can have as many as four hex digits. The actual memory address is calculated by adding a zero to the right of the segment address and adding the offset value, like this: C800:5 = C8000 + 5 = C8005. The first 640K of Windows 9x or DOS memory is called conventional memory. Look at how that memory is addressed, first in decimal and then in hex (assuming 1 kilobyte = 1024 bytes): 640K = 640 × 1024 = 655,360 There are 655,360 memory addresses in conventional memory, where each memory address can hold 1 byte, or 8 bits, of either data or program instructions. The decimal value 655,360 converted to hex is A0000 (10 × 164). So, conventional memory addresses begin with 00000h and end with A0000h minus 1h or 9FFFFh. Written in segment-and-offset form, conventional memory addresses range from 0000:0 to 9FFF:F. we know that upper memory is defined as the memory addresses from 640K to 1024K. The next address after 9FFF:F is the first address of upper memory, which is A0000, and the last address is FFFFF. Written in segment-and-offset terms, upper memory addresses range from A000:0 to FFFF:F. Here is one way to organize the conversion of a large hex value such as FFFFF to decimal (remember F in hex equals 15 in decimal).

FFFFF converted to decimal: 15 × 160 = 15 × 1 = 15 15 × 161 = 15 × 16 = 240 15 × 162 = 15 × 256 = 3,840 15 × 163 = 15 × 4096 = 61,440 15 × 164 = 15 × 65,536 = 983,040 TOTAL = 1,048,575

Remember that FFFFF is the last memory address in upper memory. The very next memory address is the first address of extended memory, which is defined as memory above 1 MB. If you add 1 to the number above, you get 1,048,576, which is equal to 1024 × 1024, which is the definition of 1 megabyte.